# Neuro Symbolic Knowledge Reasoning for Procedural Video Question Answering with Knowledge Module Learning

### v3

@@ 5.2 Analysis and Discussion
KML Ablation.
We train KML modules from scratch using only QA training, with KML-F-CLIP achieving 59.3% mean accuracy—highlighting the value of training on PKG data. KML allows exploration of multiple programs per question (see supplementary), though gains over single-program use are marginal. It also supports expert program editing for improved reliability. Evaluating different LLMs for program generation in KML-F-CLIP, GPT-4o leads with 71.6% accuracy, followed by LLaMA-3-8B (68.4), DeepSeek-V2.5 (66.5), Mistral-7B (66.2), and Qwen-2.5 (63.7), showing KML’s robustness across LLM models.
@@
Table 3: Three-hop Reasoning using KML-F-CLIP: The step-by-step reasoning outputs of KML with estimated probability value over the domain of relation using embeddings.
@@
Table 4: Four-hop Reasoning using KML-F-CLIP: The step-by-step reasoning outputs of KML with estimated probability value over the domain of relation using embeddings.
6 Conclusion
We introduced PKR-QA, a benchmark for procedural knowledge reasoning that challenges models to answer procedural questions. Our evaluation of VLMs revealed their inherent capability to leverage procedural knowledge. However, their closed-loop reasoning mechanisms lack the transparency, controllability, and domain-specific constraints required for mission-critical applications such as medical procedures or industrial automation. To address this gap, we proposed the Knowledge Module Learning (KML) framework, which explicitly grounds reasoning in procedural knowledge graphs and decouples hypothesis generation (via LLM-based program synthesis) from structured execution (via learnable knowledge modules). By constraining reasoning to predefined operations aligned with procedural logic, KML ensures reliability and interpretability without sacrificing performance. Our experiments highlight the promise of such neurosymbolic architectures, where lightweight, domain-aware modules (trained directly on KG relations) performs well in knowledge-intensive reasoning.
Acknowledgments. This research/project is supported by the National Research Foundation, Singapore, under its NRF Fellowship (Award# NRF-NRFF14-2022-0001). This research is also supported by funding allocation to Basura Fernando by the Agency for Science, Technology and Research (A*STAR) under its SERC Central Research Fund (CRF), as well as its Centre for Frontier AI Research (CFAR).

### v5

@@ VII-F Analysis and Discussion
VII-I KML Ablation.
We further study the impact of various components of our procedural knowlde question answering model shown in Figure 2. First, we ablated task/step grounding using VLM (P.VRL). Without grounding, performance dropped to 44.0%/42.4%. Then we replace the LLM-generated program with programs generated exhaustively constraining the program length to maximum of 3-hops. Given a question, we select the program using the inference mechanism of KML. Replacing LLM-generated programs with exhaustive programs yielded 57.9%/47.7%. These results show that grounding, program generation, and knowledge modules each contribute substantially to the performance. We train KML modules from scratch using only QA training, with KML-F-CLIP achieving 59.3% mean accuracy—highlighting the value of training on PKG data. KML allows exploration of multiple programs per question, though gains over single-program use are marginal. It also supports expert program editing for improved reliability.
Method Grounding (P.VRL) LLM-Gen. Program KM Use KG Accuracy Mean Accuracy
[…]
TABLE III: Ablation results showing the impact of KG training, grounding, LLM program generation, and knowledge modules.
VII-J Impact of the type of LLM for program generation
We evaluate different LLMs for program generation using KML-F-CLIP in Table IV. Most modern LLMs are good at generating the program given the question and extracting the grounded entity. We choose GPT-4o as it performs the best.
LLM acc m.acc
GPT-4o 74.6 71.6
LLaMA-3-8B 74.1 68.4
DeepSeek-V2.5 73.4 66.5
Mistral-7B 73.2 66.2
Qwen-2.5 72.6 63.7
TABLE IV: Impact of LLMs for program generation.
VII-K Hyper-Parameter Learning
We evaluate the impact of the hidden size of our knowledge modules. As shown in Figure 5, a small hidden size of 128 is sufficient to learn all modules in the KG, using a total number of 2,107,393 parameters.
Figure 5: Evaluating the impact of hidden dimension size (left) and the activation function (right) of the Knowledge modules on the validation set.
We also experimented with MSE loss to train the modules, however, the performance was poorer and the combination of contrastive loss and MSE loss did not perform as well as classical contrastive loss. The Tanh activation performs the best among all compared activation functions perhaps due to bounded Lipschitz constant of 1. Sigmoid is more contractive with Lipschitz constant of $\frac{1}{4}$ which can help stability but contributes to vanishing gradients. The grounding entity $E_{g}$ of Equation 3 is 100% accurately found by all LLM models.
VII-L KML’s logical operator performance
We now extend KML to support logical operators such as $AND$, OROR, and $NOT$ over knowledge graph relations. Consider a binary relation $r_{k}(\cdot,\cdot)$ and two head entities $e_{i}$ and $e_{j}$. Our goal is to infer the set of entities $\mathcal{E}_{ijk}$ that satisfy the logical conjunction $r_{k}(e_{i},\cdot)\land r_{k}(e_{j},\cdot)$, i.e., entities that are simultaneously related to both $e_{i}$ and $e_{j}$ under relation $r_{k}$.
Formally, this set is defined as
$\mathcal{E}_{ijk}=\bigl\{e\in\mathcal{E}\;\big|\;r_{k}(e_{i},e)\;\land\;r_{k}(e_{j},e)\bigr\}.$
We compute entity-level scores and perform logical operations directly in the score space. Let $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ denote the matrix of all entity embeddings, and let $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$ denote the output embeddings produced by the knowledge module $\phi_{R_{k}}$ for entities $e_{i}$ and $e_{j}$, respectively. Entity scores under relation $r_{k}$ are obtained via sigmoid activated similarities. Next we implement logical operators as follows:
Logical AND. The score of entities satisfying the logical $AND$ is defined as
$\text{score}_{AND}(\mathcal{E}_{ijk})=\sigma\!\left(\mathbf{E}\mathbf{e}_{i}^{\top}\right)\;\circ\;\sigma\!\left(\mathbf{E}\mathbf{e}_{j}^{\top}\right),$ (24)
where $\circ$ denotes element-wise multiplication and $\sigma(\cdot)$ is the sigmoid function. This operation corresponds to a soft intersection, assigning high scores only to entities that score highly for both relations.
Logical OR. The logical $OR$ operator is implemented as
$\text{score}_{OR}(\mathcal{E}_{ijk})=\sigma\!\left(\mathbf{E}\mathbf{e}_{i}^{\top}\right)+\sigma\!\left(\mathbf{E}\mathbf{e}_{j}^{\top}\right),$ (25)
which corresponds to a soft union, favoring entities that satisfy at least one of the relations.
Logical NOT. Finally, the logical $NOT$ operator is implemented as
$\text{score}_{NOT}\bigl(r_{k}(e_{i},\cdot)\bigr)=1-\sigma\!\left(\mathbf{E}\mathbf{e}_{i}^{\top}\right),$ (26)
which assigns high scores to entities that are unlikely to be related to $e_{i}$ under relation $r_{k}$.
To validate this, we conducted an experiment on learned knowledge modules of each relation type. We report precision at 10 performance for each module in Table V. As can be seen from the results, the performance of NOT and AND operator for most of the relations is good while the performance of OR operator is not that significant. The reason may be the logical OR operator will generate a large number of candidates for each relation and the simple scoring mechanism may not be that effective in this context. Perhaps, to implement logical $OR$ we may need to learn explicit vector function forms.
Knowledge Module NOT AND OR
HAS_GROUNDED_TOOL 100.0 94.4 24.6
HAS_PURPOSE 100.0 91.0 27.6
HAS_NEXT_STEP 100.0 60.0 82.5
HAS_TASK 100.0 70.9
HAS_STEP 100.0 97.4 37.9
HAS_ACTION 100.0 100.0 45.1
HAS_OBJECT 100.0 99.8 22.3
HAS_SIMILAR_PURPOSE 100.0 65.9 13.9
TABLE V: Precision at 10 accuracy (%) of individual Knowledge Modules under different logical operators.
VII-M KML’s Interpretability.
One key advantage of KML is its interpretability, as illustrated in the qualitative examples in Table VI and Table VII. We observe step-by-step reasoning and intermediate interpretations from the learned embeddings, offering insight into the model’s decision process. The output entities represented by the output vectors of each KM seems reasonable and accurate for the given task.
@@ VII-L KML’s logical operator performance
TABLE VI: Three-hop Reasoning using KML-F-CLIP: The step-by-step reasoning outputs of KML with estimated probability value over the domain of relation using embeddings.
@@
TABLE VII: Four-hop Reasoning using KML-F-CLIP: The step-by-step reasoning outputs of KML with estimated probability value over the domain of relation using embeddings.
[Uncaptioned image] Question: What tool could be used in the next step?
Task: Use Rice-Cooker To Cook Rice
Step: Take out some rice.
NEXT_STEP HAS_TOOL
Output=Next Step Output=Tool
cook the rice by rice cooker (0.331) rice cooker (0.233)
put the washed rice into the rice cooker (0.311) water (0.201)
soak and wash the rice (0.271) bowl (0.197)
take out some rice (0.154) water container (0.174)
pour the noodles into the water and stir (0.138) water basin (0.167)
TABLE VIII: Step-by-step reasoning of our KML method for a given question. The program [NEXT_STEP, HAS_TOOL] is executed by corresponding Knowledge Modules, producing intermediate embeddings with interpretable semantic meaning. This enables explanation of the final output. Experts can modify the program to explore alternative reasoning paths. The program is generated by an LLM based on the question and KG schema.
VIII Conclusion
We presented an extended formulation of Knowledge Module Learning (KML) as a principled neuro-symbolic framework for procedural knowledge reasoning. KML decomposes reasoning into relation-specific neural modules and composes them through explicit programmatic traversals, enabling interpretable and modular reasoning over structured procedural knowledge graphs. In contrast to black-box vision–language models, KML exposes intermediate reasoning states while retaining the generalization capability of neural representations. This journal extension introduced a formal treatment of logical operators within KML. We showed how logical reasoning over relations can be implemented as differentiable operations in score space, providing a soft relaxation of Boolean set operations compatible with gradient-based learning. We further provided a deterministic error bound for multi-hop traversal execution, offering theoretical insight into error propagation and stability under composition of learned knowledge modules. Experimental results demonstrate that KML consistently outperforms strong baselines on procedural reasoning tasks while producing interpretable execution traces aligned with relational semantics. These results, together with the theoretical analysis, establish KML as a robust foundation for structured and reliable procedural reasoning.
KML naturally supports extensions to richer logical forms, such as nested expressions, quantification, and temporal reasoning. An important future direction is to integrate KML with embodied reasoning, where procedural knowledge must be grounded in perception and action. Extending KML to embodied agents would enable reasoning over affordances, state transitions, and action outcomes, bridging symbolic procedural knowledge and real-world interaction.
Acknowledgments. This research/project is supported by the National Research Foundation, Singapore, under its NRF Fellowship (Award# NRF-NRFF14-2022-0001).